An electrical firm manufactures light bulbs that have a length of life that is normally distributed with mean equal to 809 hours and a standard deviation of 48 hours. Find the probability that a bulb burns between 728 and 848 hours


1
Expert's answer
2022-02-07T15:21:56-0500

Let "X=" a length of life: "X\sim N(\mu, \sigma^2)."

Given "\mu=809\ h, \sigma=48\ h."


"P(728<X<848)=P(X<848)-P(X\le 728)"

"=P(Z<\dfrac{848-809}{48})-P(Z\le \dfrac{728-809}{48})"

"= P(Z<0.8125)-P(X\le -1.6875)"

"\approx0.79175-0.04575= 0.7460"